OpenStudy (anonymous):

Anyone i need your help again with this problem Problem: Find the equation of the perpendicular bisector of the line joining the points A and B in the following exercises 1. A(-1,-3), B(3,5) 2. The points A(-2,3), B(6,-5), and C(8,5) are vertices of a triangle. Find the equation of the median.

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

you need to use point slope formula

OpenStudy (anonymous):

(y2-y1)=m(x2-x1) this will help you find the slope=m

OpenStudy (anonymous):

(5-(-3))=m(3-(-1)) =8=m4 =m=2 Right?

OpenStudy (anonymous):

yes i am here, thankyou for helping me hix212

OpenStudy (anonymous):

now use one of the ordered pairs and input the pair as well as the slope of 2 (y-(-3))=2(x-(-1)) y+3=2x+2 y=2x-1 now it is asking for perpendicular, so that means take the reciprocal of the slope of 2 which equals 1/2

OpenStudy (anonymous):

Do you understand this problem?

OpenStudy (anonymous):

not really undertand the last one problem. and can i ask if what formula i use in this problem Find the equation of the line passing through the points A and B in each of the following exercises 1. A (5,1), B(1,4) 2. A(-2,3) B(7,4)

OpenStudy (anonymous):

you would use point slope formula to find the m which is the slope then you will use one of the coordinates and solve for y. it will then be in y=mx+b form. try it with both points and you will find they both are equal

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

i will try it if still i cant get can i ask your help again? ehmm, i want to know also what formula use in this problem Find the equation of the perpendicular bisector of the line joining the points A and B in the following exercises 1. A(-1,-3), B(3,5) 2. The points A(-2,3), B(6,-5), and C(8,5) are vertices of a triangle. Find the equation of the median. this problem i cant really understand, so very dificult

OpenStudy (anonymous):

I did the first one for you already, but I need to look more at the second.

OpenStudy (anonymous):

for number 2, there are three median lines,so I am not sure which one to solve for.

OpenStudy (anonymous):

its okay, thankyou for helping me hix212. i will try to ask other about this problem its okay?